GNU Emacs Reference Card

(for version 21)

Starting Emacs

To enter GNU Emacs 21, just type its name: emacs

To read in a file to edit, see Files, below.

Leaving Emacs

suspend Emacs (or iconify it under X) C-z exit Emacs permanently C-x C-c

Files

read a file into Emacs C-x C-f save a file back to disk C-x C-s save all files C-x s insert contents of another file into this buffer C-x i replace this file with the file you really want C-x C-v write buffer to a specified file C-x C-w version control checkin/checkout C-x C-q

Getting Help

The help system is simple. Type C-h (or F1) and follow the directions. If you are a first-time user, type C-h t for a tutorial.

remove help window C-x 1 scroll help window C-M-v

apropos: show commands matching a string C-h a show the function a key runs C-h c describe a function C-h f get mode-specific information C-h m

Error Recovery

abort partially typed or executing command C-g recover a file lost by a system crash M-x recover-file undo an unwanted change C-x u or C-_ restore a buffer to its original contents M-x revert-buffer redraw garbaged screen C-l

Incremental Search

search forward C-s search backward C-r regular expression search C-M-s reverse regular expression search C-M-r

select previous search string M-p select next later search string M-n exit incremental search RET undo effect of last character DEL abort current search C-g

Use C-s or C-r again to repeat the search in either direction. If Emacs is still searching, C-g cancels only the part not done.

\begin{matrix}c \\ ◯ \end{matrix} 1997 Free Software Foundation, Inc. Permissions on back. v2.2

Motion

entity to move over  backward forward
character  C-b        C-f

word  M-b        M-f
line  C-p        C-n

go to line beginning (or end)  C-a        C-e
sentence  M-a        M-e
paragraph  M-{        M-}

page  C-x [      C-x ]
sexp  C-M-b      C-M-f
function  C-M-a      C-M-e

go to buffer beginning (or end)  M-<        M->

scroll to next screen C-v scroll to previous screen M-v scroll left C-x < scroll right C-x > scroll current line to center of screen C-u C-l

Killing and Deleting

entity to kill  backward forward
character (delete, not kill)  DEL        C-d

word  M-DEL      M-d

line (to end of)  M-0 C-k    C-k

sentence  C-x DEL    M-k
sexp  M-- C-M-k  C-M-k

kill region C-w copy region to kill ring M-w kill through next occurrence of char M-z char

yank back last thing killed C-y replace last yank with previous kill M-y

Marking

set mark here C-@ or C-SPC exchange point and mark C-x C-x

set mark arg words away M-@ mark paragraph M-h mark page C-x C-p mark sexp C-M-@ mark function C-M-h mark entire buffer C-x h

Query Replace

interactively replace a text string M-%

using regular expressions  M-x query-replace-regexp

Valid responses in query-replace mode are

replace this one, go on to next SPC replace this one, don’t move , skip to next without replacing DEL replace all remaining matches ! back up to the previous match ^ exit query-replace RET enter recursive edit (C-M-c to exit) C-r

Multiple Windows

When two commands are shown, the second is for “other frame.” delete all other windows C-x 1

split window, above and below  C-x 2      C-x 5 2
delete this window  C-x 0      C-x 5 0

split window, side by side C-x 3 scroll other window C-M-v

switch cursor to another window  C-x o      C-x 5 o
select buffer in other window  C-x 4 b    C-x 5 b
display buffer in other window  C-x 4 C-o  C-x 5 C-o
find file in other window  C-x 4 f    C-x 5 f
find file read-only in other window  C-x 4 r    C-x 5 r
run Dired in other window  C-x 4 d    C-x 5 d
find tag in other window  C-x 4 .    C-x 5 .

grow window taller C-x ^ shrink window narrower C-x { grow window wider C-x }

Formatting

indent current line (mode-dependent) TAB indent region (mode-dependent) C-M-\ indent sexp (mode-dependent) C-M-q indent region rigidly arg columns C-x TAB insert newline after point C-o move rest of line vertically down C-M-o delete blank lines around point C-x C-o join line with previous (with arg, next) M-^ delete all white space around point M-\ put exactly one space at point M-SPC fill paragraph M-q set fill column C-x f set prefix each line starts with C-x . set face M-g

Case Change

uppercase word M-u lowercase word M-l capitalize word M-c uppercase region C-x C-u lowercase region C-x C-l

The Minibuffer

The following keys are defined in the minibuffer. complete as much as possible TAB complete up to one word SPC complete and execute RET show possible completions ? fetch previous minibuffer input M-p fetch later minibuffer input or default M-n regexp search backward through history M-r regexp search forward through history M-s abort command C-g Type C-x ESC ESC to edit and repeat the last command that used the minibuffer. Type F10 to activate the menu bar using the minibuffer.

GNU Emacs Reference Card

Buffers

select another buffer C-x b list all buffers C-x C-b kill a buffer C-x k

Transposing

transpose characters C-t transpose words M-t transpose lines C-x C-t transpose sexps C-M-t

Spelling Check

check spelling of current word M-$ check spelling of all words in region M-x ispell-region check spelling of entire buffer M-x ispell-buffer

Tags

find a tag (a definition) M-. find next occurrence of tag C-u M-.

specify a new tags file  M-x visit-tags-table

regexp search on all files in tags table M-x tags-search run query-replace on all the files M-x tags-query-replace continue last tags search or query-replace M-,

Shells

execute a shell command M-! run a shell command on the region M-| filter region through a shell command C-u M-| start a shell in window *shell* M-x shell

Rectangles

copy rectangle to register C-x r r kill rectangle C-x r k yank rectangle C-x r y open rectangle, shifting text right C-x r o blank out rectangle C-x r c prefix each line with a string C-x r t

Abbrevs

add global abbrev C-x a g add mode-local abbrev C-x a l add global expansion for this abbrev C-x a i g add mode-local expansion for this abbrev C-x a i l explicitly expand abbrev C-x a e

expand previous word dynamically M-/

Regular Expressions

any single character except a newline . (dot) zero or more repeats * one or more repeats + zero or one repeat ? quote regular expression special character c \c alternative (“or”) \| grouping \( … \) same text as nth group \n at word break \b not at word break \B

entity  match start match end
line  ^$

word  \<           \>

buffer  \`           \'
class of characters  match these match others
explicit set  [ …]       [^…]
word-syntax character  \w           \W
character with syntax c  \sc  \Sc

International Character Sets

specify principal language  M-x set-language-environment
show all input methods  M-x list-input-methods

enable or disable input method C-\ set coding system for next command C-x RET c

show all coding systems  M-x list-coding-systems

choose preferred coding system M-x prefer-coding-system

Info

enter the Info documentation reader C-h i find specified function or variable in Info C-h C-i

Moving within a node:

scroll forward SPC scroll reverse DEL beginning of node . (dot)

Moving between nodes:

next node n previous node p move up u select menu item by name m select nth menu item by number (1–9) n follow cross reference (return with l) f return to last node you saw l return to directory node d go to any node by name g

Other:

run Info tutorial h quit Info q search nodes for regexp M-s

Registers

save region in register C-x r s insert register contents into buffer C-x r i

save value of point in register C-x r SPC jump to point saved in register C-x r j

Keyboard Macros

start defining a keyboard macro C-x ( end keyboard macro definition C-x ) execute last-defined keyboard macro C-x e append to last keyboard macro C-u C-x (

name last keyboard macro  M-x name-last-kbd-macro

insert Lisp definition in buffer M-x insert-kbd-macro

Commands Dealing with Emacs Lisp

eval sexp before point C-x C-e eval current defun C-M-x

eval region  M-x eval-region

read and eval minibuffer M-: load from standard system directory M-x load-library

Simple Customization

customize variables and faces M-x customize

Making global key bindings in Emacs Lisp (examples):

(global-set-key "\C-cg" 'goto-line)
(global-set-key "\M-#" 'query-replace-regexp)

Writing Commands

(defun command-name (args)

"documentation" (interactive "template")

body)

An example:

(defun this-line-to-top-of-window (line)
  "Reposition line point is on to top of window.
With ARG, put point on line ARG."
  (interactive "P")
  (recenter (if (null line)
                0
              (prefix-numeric-value line))))

The interactive spec says how to read arguments interactively. Type C-h f interactive for more details.

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1997 Free Software Foundation, Inc.

v2.2 for GNU Emacs version 21, 1997 designed by Stephen Gildea

Permission is granted to make and distribute copies of this card provided the copyright notice and this permission notice are preserved on all copies. For copies of the GNU Emacs manual, write to the Free Software Foundation, Inc., 59 Temple Place, Suite 330, Boston, MA 02111-1307 USA